#3 only. Thanks!
take B 2.5 T, v 4 m/s, 30 cm, and 2. In the generator rod increase in magnetic flux in one 4 current 2.0 A, clockwise. Find (a) the rod-the external R 2 second, (b) the emf, (c) the magnetic back-force on the at which electrical force needed to keep it moving. Calculate the rate force does energy is produced by the emf, and which the external 1.50]. work, showing equality. (e) Find the resistance of the circuit 3. In #2, now assume that a second, motor-rod is moving at 3.2 m/s, with all given data the same. Find (a) the magnetic force on the motor-rod and the back-emf on it. (c) Find the net emf, and use this to find (d) the resistance r in this circuit. (e) Find the rate of production of mechanical energy and of heat, and check consistency with a preceding result. What determines which rod is the generator and which the motor? What would happen if the two rods moved at the same speed?
Answer #1


magnetic flux = B*A = B*l*dx


rate of change in flux = B*i*dx/dt = B*l*v = 2.5*0.3*4 = 3 V


------------


(b)


emf = rate of change in flux = 3 V


-------------

(c)


F = I*l*B = 2*0.3*2.5 = 1.5 N

-------------------


(d)


rate of electrical energy = e*I = 3*2 = 6 W

(ii)


Work = F*v = 1.5*4 = 6 W


==============


(e)


rate of energy = i^2*R

R = 6/2^2 = 1.5 ohms

============================================

(